The circumference of one wheel is 20mm longer than of  another wheel. The product of the two circumference is 144300mm. Calculate the dimension of the two wheels.

Grade:12th pass

1 Answers

Harshit Singh
3 years ago
Dear Student

given that
c1=c2+20
and
c1*c2=144300 mm^2
therefore
(c2+20)*c2=144300

hence c2^2+20*c2-144300=0
this is a quadratic equation in c2 and can be solved using the standard
procedure
as c2=(-20+760)/2= 370
The negative value will be ommitted.
c1= c2+20
c1=390.

circumference= 2*pi*r
finding radius from the obtained circumference.
r1= 58.88 mm
r2= 62.07 mm

Thanks